(Problem #30, page 429) An unknown distribution has a mean of 19 and a standard deviation of 20. Let X = the object of interest. What is the sample size if the mean of ΣX is 15,200?
Answer #1

Mean = X / n

19 = 15200 / n

Solve for n

n = 15200 / 19

n = 800

Sample size = 800